Polyethylene is a semi-crystalline polymer which can adapt a metallic-like structure. It has an orthorhombic structure with a = 7.41, b = 4.94 and c =2.55 all in angstroms (10-10 m). If there are two mer units per unit cell, calculate the density of crystalline polyethylene
1
Expert's answer
2020-07-29T06:35:47-0400

orthorhombic structure with a = 7.41, b = 4.94 and c =2.55

Volume of orthorhombic V= a×b×c

=7.41×4.94×2.55 ×10-30 m

= 93.35 ×10-30 m

Density of orthorhombic

= nA/VNA

=2×28.05/93.35×10-30×6.023×1023

= 9.98×105